Hello my question is " Point Z is translated using the rule (x,y)----> (x-4, y+11). The coordinates of Z one are (4,-11). What are the coordinates of point Z?
Answer by jim_thompson5910(35256)

"Z one"? I think you mean "Z prime" or Z'

If Z' is (4,-11), then

x-4 = 4

x = 4+4

x = 8

y+11 = -11

y = -11-11

y = -22

So if Z' is (4,-11), then Z is (8, -22)

Basically, if you start with Z(8, -22) and you apply the rule (x,y)----> (x-4, y+11), then you will get Z'(4,-11)